# Graph Representation of Narrative Context: Coherence Dependency via Retrospective Questions

### v1

@@ 4 Study I: Edge Efficacy
Our first study examines the graph edges on whether they express useful relations. Ideally, a non-empty edge should bridge coherence between two nodes through its retrospective questions. For appropriate assessment, we adopt the recap identification on RECIDENT dataset Li et al. (2024), a task on narratives that identifies whether certain preceding snippets can function as a recap or prelude to the audience in regards to a current context.
Concretely, the input takes a short snippet from a novel or show script, along with a provided list of its preceding snippets; this task resolves which preceding snippets are directly related with the current one in terms of plot progression, requiring contextual understanding of narrative development. As NarCo is proposed to capture the inter-node coherence relations, edges of retrospective questions could be leveraged to link the current snippet to related preceding ones. Therefore, RECIDENT serves as a natural testbed for comprehensive evaluation of edge efficacy.
@@ 4.1 Approach
For this study, our proposed approach targets upon the zero-shot baseline with LLMs in Li et al. (2024), where ChatGPT is originally asked to select the related recap snippets from the list of preceding candidates based on their context.
With NarCo, we regard each current snippet as a target graph node $v_{t}$, and the list of its $N$ preceding snippets $\{v_{c}|c=1,..,N\}$ as the candidates. For $v_{t}$ and each of its candidate $v_{c}$, the edge is realized denoted by $e_{ct}$. As each question in $e_{ct}$ should reflect their causal or temporal relations, we utilize these questions directly from two distinct aspects.
@@
Normally, each snippet is represented by its context as in the baseline. To evaluate the coherence depicted by edges, we instead propose to represent each snippet solely based on the edge relations: for a candidate node $v_{c}$, we use its concatenated questions in $e_{ct}$ for representation, denoted by $\{q_{c}|c=1,..,N\}$, and completely neglect the original context, so to ensure an entirely isolated assessment of edge relations.
Specifically, given the context of the target snippet $v_{t}$, we now ask LLM to select which $q_{c}$ addresses important questions asking recap information significant to comprehend the current context. As each $q_{c}$ contains multiple questions, we ask LLM to score each $q_{c}$ in $[0,5]$, with higher scores indicating better overall questions. Candidates with empty edges are directly assigned 0 score.
@@
Alternatively, as mentioned in Section 3.1, the edge degree (number of questions) could suggest how cohesively related between two nodes. To this end, we further propose to simply deem the edge degree as the score for ranking candidates, without any inference on the node context or edge relations at all. Though being rather unconventional, ranking recaps by edge degrees could approximately reflect the edge quality.
For either the relation score or degree score, it could be used standalone or interpolated with the baseline selection. More formally, we obtain the rank $\in[1,N]$ of each candidate $i$ by relation scores, denoted as $r^{rel}_{i}$, and the rank by degree scores as $r^{deg}_{i}$, along with the binary selection $b_{i}$ from baseline. The final score $s$ of each candidate is:
@@
$\mathbb{I}$ is the indicator function that boosts the baseline selection by $\lambda$ rank; relation and degree ranks are interpolated by $\alpha$ and $\beta$. The final score is then ranked to select top candidates with recap information (lower is better). Setting these to 0 accordingly can thereby evaluate each method standalone.
@@ Methods
4.3 Results and Analysis
Table 1 shows the zero-shot evaluation results on the test set of RECIDENT. Notably, the interpolation with NarCo edges (Full) consistently brings significant improvement upon the baseline (BL), by 4.9 / 2.4 F1 on NDP / GOT respectively with ChatGPT, up to a 21.7% relative improvement. The stronger GPT-4 boosts performance for all methods as expected, and still advancing 3.5 / 4.7 F1 upon BL on NDP / GOT as well.
Moreover, selection solely based on edge relations without disclosing the context (Rel) could obtain comparable or better performance than the baseline, with the only exception of ChatGPT on GOT. Overall, Table 1 demonstrates the effective utility of NarCo leveraging its edge efficacy, offering a complementary enhancement.
For more in-depth insights, we further perform two additional evaluation with ChatGPT:
  • Full-F: the Full setting with all generated questions, without the back verification stage.
Corresponding results are shown in Table 2, where ranking by edge degrees of NarCo exhibits decent performance. It even surpasses the baseline on NDP by 1+%, which is impressive for the fact that it does not undergo any task-specific inference. Understandably, it indeed lags behind the baseline on GOT by a noticeable margin. As for Full-F, the trivially degraded performance suggests that, our proposed approach can be quite robust against noisy edge questions, as LLM assigns high scores as long as under the presence of good questions.
@@ 5 Study II: Node Augmentation
Our second study underscores the NarCo utility of local context augmentation, examining whether the graph typology could enrich the node representation with global contextual information.
Specifically, for a node $v_{j}$, a preceding node $v_{i}$ and succeeding node $v_{k}$ such that $i<j<k$, $e_{ij}$ depicts outgoing questions arising from $v_{j}$ to $v_{i}$, and $e_{jk}$ specifies incoming questions from $v_{k}$ that can be clarified by $e_{j}$. These questions either highlight important aspects of events or situations in the current context, or provide implication of subsequent development. Such auxiliary information from neighboring nodes is especially useful for retrieval on narratives, as each passage is not independent and rather being related with others.
We hence investigate if an embedding function on top of NarCo could lead to enriched local representation. Towards this objective, we consider the plot retrieval task defined in Xu et al. (2023b), which aims to find the most relevant story snippets given a query of short plot description. It is challenging as queries are often abstract based on readers’ overall understanding of the stories, requiring essential background information clarified on candidates, similar to the concept of decontexualization Choi et al. (2021). Retrieval on narratives thereby fits our evaluation purpose well.
@@ 5.1 Approach
For this task, candidate snippets from stories are retrieved upon a given query. We build the graph for the full narrative, e.g. an entire novel, according to Section 3 and regard all candidate snippets as graph nodes to be retrieved from. Our proposed approach focuses on fusing edge questions into the node representation for enhanced retrieval.
Xu et al. (2023b) follows the classic paradigm of contrastive learning that learns a BERT-based encoder Devlin et al. (2019) on queries and candidates. As its trained model is not released yet, our approach adopts the public BGE encoder Xiao et al. (2023) in this work that ranks top on the MTEB leaderboard [^1] https://huggingface.co/spaces/mteb/leaderboard. For comprehensive evaluation, we propose methods with NarCo for both zero-shot and supervised settings.
@@ 5.1.2 Supervised Learning
We then introduce our proposed supervised approach that reranks candidates with augmented node embeddings. Specifically, the enrichment is formulated as an attention, with the user query as query, edge questions as both key and value, such that a new embedding is obtained upon all edge questions conditioned on the query. Let $\mathcal{A}_{i}$ be the attention scores of the ith node, the augmented node embedding $\mathbf{h}^{a}_{i}$ is denoted as:
@@
With the augmented embedding for the ith node $\mathbf{h}^{a}_{i}$, the model simply reranks top retrieved candidates from a baseline system for inference. It is trained with the supervised contrastive loss Khosla et al. (2020) to maximize the similarity between each query $q$ and its positive targets $P(q)$ among $N$ in-batch candidates (details in Appx A.3):
@@ 5.2 Experiments
A query may have one to many positive snippets (up to 7). We take the typical information retrieval metric normalized Discounted Cumulative Gain (nDCG), assigning the same relevance for each positive snippet equally.
@@ Methods
  • Zero Shot (ZS): the baseline method that ranks candidates based on query-node similarity.
@@
  • Supervised (SU): the baseline model trained supervisedly on queries and candidates only.
  • SU+NarCo: our proposed rerank model that utilizes global-contextualized embeddings; the inference reranks top 50 candidates by SU.
5.3 Results
Table 3 shows the evaluation results of our settings. Notably, our proposed zero-shot interpolation with query-edge similarity improves upon its baseline on all nDCG metrics, leading 3.4% on nDCG@10 ($\lambda=0.1$), which corroborates the positive information gain from edges for direct node augmentation. The same trend still holds up for the supervised model, improving by a large margin, especially by 2.4% on nDCG@1. Overall, NarCo is shown beneficial towards the acquisition of better local embeddings, demonstrated useful for retrieval with our proposed utilization, which in turn advocates the motivation of this research to push the global context modeling among local segments that fosters a more nuanced comprehension.
@@ 5.3 Results
6 Study III: Application in QA
Our last study sheds light on graph utility in broader applications, moving beyond the focus of graph edges and nodes themselves. We choose QuALITY Pang et al. (2022), a multi-choice question answering (QA) dataset on long documents, mostly on fiction stories from Project Gutenberg. With an averaged length of 5k+ tokens per document, we investigate the potentials for retrieval-based approaches, where NarCo may assist to recognize more relevant context, leading to better QA performance benefited from enhanced retrieval.
Specifically, questions in QuALITY were constructed with global evidence in mind, demanding multiple parts in the document to reason upon. In this work, we target the zero-shot QA evaluation, leveraging NarCo to obtain more accurate context during the retrieval process.
@@ Methods
Retrieval-based approaches are commonly adopted for tackling long context. As experimented by Pang et al. (2022); Xu et al. (2024), we also split the full document by short snippets and retrieve relevant snippets with regard to the question. We apply the same retrieval process described in Section 5.1.1, where the query-edge similarity is interpolated as in Eq (2) using BGE-Large encoder. The retrieved snippets are then concatenated as the shortened relevant context for subsequent QA.
@@
We employ Llama2 Touvron et al. (2023) and ChatGPT for the zero-shot QA inference. As evaluation on the test set requires submission to the ZeroSCROLLS leaderboard Shaham et al. (2023), we first perform fine-grained analysis on the dev set with short retrieved context ($<$1k), then submit the final test results using ChatGPT with 1.5k context limit, aligned with Xu et al. (2024) for direct comparison. Baseline retrieval and our Enhanced retrieval are denoted by R and ER.
Table 4 & 5 present the evaluation results on the dev set and test set respectively. Results on the dev set suggest that ER enhanced by NarCo can boost QA performance with all LLMs, especially with the smaller 7B model by 5% accuracy, fulfilling our initiative to utilize NarCo in broader applications. The improvement from superior retrieved context is consistent, further confirmed by the 2% leading margin with ChatGPT on both dev and test set.
For this QA type of tasks, potential enhancement may happen in the retrieval process or the actual QA inference. As we have demonstrated NarCo could assist retrieval that leads to improved QA performance, the QA inference itself has not received auxiliary signals from edge questions. We leave room for future research on the integration of edges into the QA inference for further improvement.
@@
Llama2-7B 40.97 (± 0.67) 45.97 (± 0.63)
Llama2-70B 61.56 (± 0.06) 63.98 (± 0.23)
ChatGPT 63.66 (± 0.06) 65.92 (± 0.34)
@@
Table 5: Evaluation results on the test set of QuALITY submitted to the ZeroSCROLLS leaderboard. Accuracy of ChatGPT* is provided by the ZeroSCROLLS organizers; Llama2-70B (+R)* is reported by Xu et al. (2024). Settings with +R or +ER are directly comparable (all within 1.5k context limit).
@@ 7 Conclusion
We introduce NarCo, a novel paradigm of narrative representations using a graph structure composed of snippet nodes connected by their coherence dependencies. The edges are formulated as retrospective questions that find background information from prior snippets to enhance comprehension of the current snippet. To realize this concept without human annotations, we propose a two-stage LLM prompting approach to generate these questions. NarCo facilitates narrative understanding by offering informative coherence relationships between snippets explicitly and enriched snippet embeddings with global context, validated by positive results on recap identification and plot retrieval tasks. We additionally utilize NarCo in a long-context QA task to further demonstrate its practical utility in downstream applications.
@@ A.1 Full Prompts and Details
Full prompts of the two-stage LLM prompting (Section 3) are provided in Figure 3-5. We specify the maximum number of generated questions for a node pair as 4 in the prompt.
For the task of plot retrieval (Section 5) and long context QA (Section 6), we construct edges within a neighboring window of 4 preceding nodes, such that the graph realization is proportional to the input instead of being quadratic. For recap identification Li et al. (2024), edges are obtained on the provided preceding snippets.
For a context with $T$k tokens, it takes approximately $6T$k tokens to obtain all edge questions of NarCo using GPT-4, which costs \$$0.06T$ as of this writing.
@@ A.2 Qualitative Examples
Examples of generated questions on Game of Thrones from recap identification Li et al. (2024).

### v2

@@ 4 Study I: Edge Efficacy
Our first study examines the graph edges on whether they express useful relations, such that the generated retrospective questions should bridge the coherence between two context. For appropriate assessment, we conduct recap identification on RECIDENT dataset Li et al. (2024), a task on narratives that identifies whether certain preceding snippets can function as a recap or prelude to the audience in regards to a current context.
Concretely, the input takes a short snippet from a novel or show script, along with a provided list of its preceding snippets. The task resolves which preceding snippets are directly related with the current one in terms of plot progression, requiring contextual understanding of narrative development. As NarCo is designed to capture the coherence relations between context, edges of retrospective questions could be leveraged to link the current snippet to related preceding ones. Therefore, RECIDENT serves as a natural testbed for comprehensive evaluation of edge efficacy.
@@ 4.1 Approach
For this study, our proposed approach targets upon the zero-shot baseline with LLMs in Li et al. (2024), where ChatGPT is originally asked to select the related recap snippets from the list of preceding candidates based on their text content.
With NarCo, we regard each current snippet as a target graph node $v_{t}$, and the list of its $N$ preceding snippets $\{v_{c}|c=1,..,N\}$ as the candidates. For $v_{t}$ and each of its candidate $v_{c}$, the edge is realized as $e_{ct}$. As questions in $e_{ct}$ should reflect causal or temporal coherence, we directly utilize these questions in two following ways.
@@
Normally, candidate snippets are processed by their text content as in the baseline. To evaluate the coherence depicted by edges, we instead propose to identify recap snippets solely based on the edge relations: for a candidate node $v_{c}$, we concatenate all its questions in edge $e_{ct}$, denoted by $q_{c}$, to identify recap, and completely neglect original text content, so to ensure an entirely isolated assessment of edge relations.
Specifically, given the context of a target snippet $v_{t}$, and $N$ candidates $\{q_{c}|c=1,..,N\}$ represented by questions, we now ask a LLM to score which $q_{c}$ addresses important questions that are significant to comprehend the current context, with higher scores indicating better overall questions that provide recap information. Candidates with empty edges are directly assigned 0 score.
@@
Alternatively, as mentioned in Section 3.1, the number of questions between two nodes could suggest how cohesively related they are. We take this number as the edge degree, and propose to simply deem it as the score to rank candidates, without any inference on the node context or edge relations at all. Though being rather unconventional, ranking candidates by edge degrees further reflects the edge quality.
With either the relation score or degree score, it can be used standalone or interpolated with the baseline selection. More formally, we obtain the rank $\in[1,N]$ of each candidate $i$ by relation scores, denoted as $r^{rel}_{i}$, and the rank by degree scores $r^{deg}_{i}$, along with binary selection $b_{i}$ from the original baseline. The final score $s$ of each candidate is:
@@
$\mathbb{I}$ is the indicator function that boosts the baseline decision $b_{i}$ by $\lambda$ rank; relation and degree ranks are interpolated by $\alpha$ and $\beta$. The final score is then ranked to select top candidates with recap information (lower is better). Setting $\alpha$/$\beta$/$\lambda$ to 0 can thereby evaluate each method standalone.
@@ Methods
4.3 Results
Table 1 shows the zero-shot evaluation results on the test set of RECIDENT. Notably, the interpolation with NarCo edges (Full) consistently brings improvement upon the baseline (BL), by 4.9 / 2.4 F1 on NDP / GOT respectively with ChatGPT, up to a 21.7% relative improvement. The stronger GPT-4 boosts performance for all methods as expected, and NarCo still advances 3.5 / 4.7 F1 upon BL on NDP / GOT as well.
Moreover, selection solely based on edge relations (Rel) obtains comparable or better performance than the baseline, with the only exception on GOT with ChatGPT. Overall, Table 1 effectively demonstrates the edge efficacy of NarCo that expresses coherence through retrospective questions.
For in-depth analysis, we further perform two additional evaluation with ChatGPT:
  • Full-F: the Full setting with all generated questions, without Filtering by self verification.
@@ 4.3 Results
Table 2 shows the additional evaluation results, where ranking by edge degrees of NarCo exhibits decent performance. It even surpasses the baseline on NDP by 1+%, which is impressive for the fact that it does not undergo any task-specific inference. Understandably, it indeed lags behind the baseline on GOT by a noticeable margin.
For Full-F, the degradation is trivial from Full. It is also expected, as the LLM scoring on relations is based on the presence of “good” questions that reflect recap information, which should be retained by the verification stage. Thus, our approach with NarCo is shown robust against noisy questions.
4.4 Graph Insights
The majority of generated questions in NarCo are what/why/how-type of questions. Their ratios are provided in Table 3, along with the averaged number of questions per edge before / after the self verification stage (Section 3.2).
NDP GOT
What-Questions Ratio 61.5% 58.4%
Why-Questions Ratio 26.5% 25.2%
How-Questions Ratio 7.8% 14.0%
# Questions per Edge 3.4 3.5
+ Self Verification 1.9 2.0
Table 3: Statistics of NarCo in Study I (Section 4).
@@ 5 Study II: Node Augmentation
Our second study underscores the NarCo utility of local context augmentation, examining whether the graph typology could enrich node representation with global contextual information.
Specifically, for a node $v_{j}$, a preceding node $v_{i}$ and succeeding node $v_{k}$ such that $i<j<k$, $e_{ij}$ depicts outgoing questions arising from $v_{j}$ to $v_{i}$, and $e_{jk}$ specifies incoming questions from $v_{k}$ that can be clarified by $e_{j}$. These questions either highlight important aspects of events or situations in the current context, or provide implication of subsequent development. Such auxiliary information from neighboring nodes is especially useful for retrieval on narratives, as each passage tends to be more interconnected with others than isolated.
We hence investigate if an embedding function on top of NarCo could lead to enriched local representation. Towards this objective, we consider the plot retrieval task defined in Xu et al. (2023b), which aims to find the most relevant story snippets given a query of short plot description. It is challenging as queries are often abstract based on readers’ overall understanding of the stories, requiring essential background information clarified on candidates, analogous to the concept of decontexualization Choi et al. (2021). Retrieval on narratives thereby fits our evaluation purpose well.
@@ 5.1 Approach
For this task, candidate snippets from stories are retrieved upon a given query. We regard all candidate snippets as graph nodes to be retrieved from, and derive NarCo edges of neighboring nodes. Our proposed method focuses on fusing edge questions into node representation for enhanced retrieval.
Xu et al. (2023b) follows the classic paradigm of contrastive learning that learns a BERT-based encoder Devlin et al. (2019) on queries and candidates. As its trained model is not released as of this writing, our approach adopts the public BGE encoder Xiao et al. (2023) in this work that ranks top on the MTEB leaderboard [^1] https://huggingface.co/spaces/mteb/leaderboard. For comprehensive evaluation, we propose methods with NarCo for both zero-shot and supervised settings.
@@ 5.1.2 Supervised Learning
We then introduce our proposed supervised approach that reranks candidates with augmented node embeddings. Specifically, the enrichment is formulated as an attention, with the user query as query, edge questions as both key and value, such that a new node embedding is obtained attending its edge questions conditioned on the query. Let $\mathcal{A}_{i}$ be the attention scores of the ith candidate node, the augmented node embedding $\mathbf{h}^{a}_{i}$ is denoted as:
@@
With the augmented embedding for the ith node $\mathbf{h}^{a}_{i}$, the model simply reranks top retrieved candidates from a baseline system. It is trained with the supervised contrastive loss Khosla et al. (2020) to maximize the similarity between each query $q$ and its positive targets $P(q)$ among $N$ in-batch candidates (details in Appx A.3):
@@ 5.2 Experiments
A query may have one or many positive snippets (up to 7). We take the typical information retrieval metric normalized Discounted Cumulative Gain (nDCG), assigning the same relevance for each positive snippet equally.
@@ Methods
  • Zero Shot (ZS): the zero-shot method that ranks candidates based on the query-node similarity.
@@
  • Supervised (SU): the baseline supervised model without leveraging NarCo.
  • SU+NarCo: our proposed rerank model that utilizes the global-contextualized embeddings; the inference reranks top 50 candidates by SU.
@@
5.3 Results
Table 4 shows the evaluation results of the four settings. Notably, our proposed zero-shot interpolation with query-edge similarity improves upon its baseline on all nDCG metrics, leading 3.4% on nDCG@10 ($\lambda=0.1$), confirming the positive information gain from edges that contribute useful contextual information. The same trend still holds up for the supervised model that learns enriched embeddings leveraging edge relations, especially by the 2.4% improvement on nDCG@1.
Overall, NarCo is shown helpful towards the acquisition of better local representation, through the explicit relational dependencies beyond local context. The empirical results advocate the direction of fine-grained context modeling, which could foster a more nuanced comprehension.
6 Study III: Broader Application
Our last study sheds light on the potentials of graph utility in broader applications. As a first step towards this new direction, in this work, we evaluate with Retrieval-Augmented Generation (RAG) Lewis et al. (2020) in the task of long document question answering. Experiments are conducted on QuALITY Pang et al. (2022), a multi-choice QA dataset on narrative documents, mostly being fiction stories from Project Gutenberg. With an averaged length of 5k+ tokens per document, we adopt the retrieval-based approaches, where relevant snippets conditioned on the question are retrieved first, then fed to a LLM to generate answers, following a standard RAG paradigm.
Especially, QuALITY was constructed with global evidences in mind: questions may require multiple parts in the document to reason upon. Therefore, NarCo may assist to recognize more relevant snippets through the extracted relations across the narrative context, leading to improved QA performance benefited from enhanced retrieval.
@@
Retrieval-based approaches are commonly adopted for tackling long context, which have been evaluated on QuALITY by previous works Pang et al. (2022); Xu et al. (2024); Sarthi et al. (2024). Following these setup, we split the full document by short snippets, and retrieve relevant snippets with regard to the question, which are then concatenated as the shortened context for subsequent zero-shot QA inference by LLMs. To leverage NarCo, we apply the same retrieval process described in Section 5.1.1 to identify relevant snippets, where the query-edge similarity is interpolated as in Eq (2) using BGE-Large encoder.
@@
We employ Llama2 Touvron et al. (2023) and ChatGPT for the zero-shot QA inference. As evaluation on the test set requires submission to the ZeroSCROLLS leaderboard Shaham et al. (2023), we first perform fine-grained performance analysis on the dev set with short retrieved context ($<$1k tokens), then submit the final test set results using ChatGPT with 1.5k context limit, aligned with Xu et al. (2024) for direct comparison. The baseline retrieval method and our Enhanced retrieval are denoted by R and ER respectively.
Table 5 & 6 present the evaluation results on the dev set and test set respectively. Results on the dev set suggest that ER can boost QA performance with all LLMs, especially with the smaller 7B model by 5% accuracy, fulfilling our initiative to effectively utilize NarCo in broader applications. The improvement from enhanced context retrieval is consistent, further confirmed by the 2% leading margin with ChatGPT on both the dev and test set.
Having demonstrated that NarCo can improve RAG in narratives through enhanced retrieval, its utility beyond the retrieval process may be further exploited, e.g. potential facilitation on LLM pretraining or inference directly. We leave future research to explore additional integration of fine-grained context modeling.
@@
Llama2-7B 40.97 (± 0.67) $\longrightarrow$ 45.97 (± 0.63)
Llama2-70B 61.56 (± 0.06) $\longrightarrow$ 63.98 (± 0.23)
ChatGPT 63.66 (± 0.06) $\longrightarrow$ 65.92 (± 0.34)
@@
Table 6: Evaluation results on the test set of QuALITY submitted to the ZeroSCROLLS leaderboard. Accuracy of ChatGPT* is provided by the ZeroSCROLLS organizers; Llama2-70B (R)* is reported by Xu et al. (2024). Performance of three retrieval-based experiments are directly comparable (same 1.5k context limit). We exclude another related work RAPTOR Sarthi et al. (2024), as they use smaller QA models and different context limit, thus not directly comparable.
@@ 7 Conclusion
We address the distinctive characteristics of narratives, and propose a novel paradigm of fine-grained context modeling, which explicitly captures the inter-connective coherence within narrative context. A graph is thereby formulated, dubbed NarCo, with edges encompassing free-form retrospective questions to depict the relational dependencies. NarCo is practically realized by LLMs via our designed two-stage prompting scheme, leveraging the promising development of LLMs without reliance on human annotations. To examine the graph properties and its utility, three unique studies are conducted, where NarCo is shown to bring empirical improvement on various narrative applications.
@@ A.1 Full Prompts and Details
Full prompts of our designed two-stage LLM prompting scheme (Section 3) are provided in Figure 3-5. We specify the maximum number of generated questions for a node pair as 4 in the prompt.
For the task of plot retrieval (Section 5) and long document QA (Section 6), we construct edges within a neighboring window of 4 preceding nodes, such that the graph realization is proportional to the input instead of being quadratic. For recap identification Li et al. (2024), edges are obtained on the provided preceding snippets.
For a context with $T$k tokens, it takes approximately $6T$k tokens to obtain all edge questions of NarCo using GPT-4, which costs \$$0.03T$ as of this writing.
@@ A.2 Qualitative Examples
Examples of generated questions on Game of Thrones from RECIDENT Li et al. (2024).